### v1

# Covering Complete Geometric Graphs by Monotone Paths

Abstract
Given a set $A$ of $n$ points (vertices) in general position in the plane, the complete geometric graph $K_{n}[A]$ consists of all (${n\choose 2}$ segments (edges) between the elements of $A$. It is known that the edge set of every complete geometric graph on $n$ vertices can be partitioned into $O(n^{3/2})$ noncrossing paths (or matchings). We strengthen this result under various additional assumptions on the point set. In particular, we prove that for a set $A$ of $n$ randomly selected points, uniformly distributed in $[0,1]^{2}$, with probability tending to $1$ as $n\rightarrow\infty$, the edge set of $K_{n}[A]$ can be covered by $O(n^{4/3}t(n))$ noncrossing paths (or matchings), where $t(n)$ is any function tending to $\infty$. On the other hand, we construct $n$-element point sets such that covering the edge set of $K_{n}(A)$ requires a quadratic number of monotone paths.
keywords:
convexity, geometric graph, complete graph, crossing family, plane subgraph
1 Introduction
A set of points in the plane is said to be (i) in general position if no $3$ points are collinear; and (ii) in convex position if every point in the set is an extreme point of the convex hull. Given $n$ points in general position in the plane, the graph obtained by connecting certain point-pairs by straight-line segments is called a geometric graph $G$. If no two segments (edges) of $G$ cross each other, then $G$ is said to be noncrossing or a plane graph.
In $2005$, Araujo, Dumitrescu, Hurtado, Noy, and Urrutia [3] asked the following question: Can the edge set of every complete geometric graph on $n$ vertices be partitioned into a small number of noncrossing matchings? In particular, they asked:
[…]
Does there exist a constant $c>0$ such that every complete geometric graph on $n$ vertices can be partitioned into at most $cn$ plane matchings?
It is easy to verify that when $n$ is even (resp., odd), every complete geometric graph of $n$ vertices in convex position can be decomposed into $n-1$ (resp., $n$) perfect plane matchings. On the other hand, the best known upper bound for point sets in general position is only $O(n^{3/2})$ [3], which seems to be far off.
In this paper we study a variant of the above question, namely the problem of covering complete geometric graphs by monotone paths, a special type of plane paths. A path (or matching) is called monotone if there is a straight line $\ell$ such that no line orthogonal to $\ell$ intersects it in more than one point. How many paths does one need in this case? Note that every path naturally decomposes into two matchings.
Monotone paths are ubiquitous and have been studied extensively over the years, particularly in the field of optimization [6]. On the other hand, noncrossing paths and matchings have been an attraction for geometers [8, 9, 10]. Moreover, every monotone path is noncrossing.
Before stating our results, we clarify the terminology.
Definitions and notations.
For a triple of points $a,b,c$, let $\Delta{abc}$ denote the triangle with vertices $a,b,c$. All point sets appearing in this paper are assumed to be in general position, and the logarithms are in base $2$.
A polygonal path $\xi=(v_{1},v_{2},\ldots,v_{t})$ in $\mathbb{R}^{2}$ is monotone in direction $\mathbf{u}\in\mathbb{R}^{2}\setminus\{\mathbf{0}\}$ (or is $\mathbf{u}$-monotone) if every directed edge of $\xi$ has a positive inner product with $\mathbf{u}$, that is, $\langle\overrightarrow{v_{i}v_{i+1}},\mathbf{u}\rangle>0$ for $i=1,\ldots,t-1$; here $\mathbf{0}$ is the origin. A path $\xi=(v_{1},v_{2},\ldots,v_{t})$ is monotone if it is monotone in some direction $\mathbf{u}\in\mathbb{R}^{2}\setminus\{\mathbf{0}\}$. Obviously, any monotone path is noncrossing.
A matching is said to be monotone in direction $\mathbf{u}\in\mathbb{R}^{2}\setminus\{\mathbf{0}\}$ if the edge projections onto $\mathbf{u}$ are disjoint intervals. A matching is said to be monotone if it is monotone in some direction $\mathbf{u}\in\mathbb{R}^{2}\setminus\{\mathbf{0}\}$. Equivalently, a matching is monotone if its edge set is the subset of the edge set of a monotone path.
For any constant $\alpha>0$, a set of $n$ points is $\alpha$-dense if the ratio between the longest and the shortest distance between two points in the set is at most $\alpha\sqrt{n}$.
Our results.
Given a point-set $A$, let $K_{n}[A]$ denote the complete geometric graph induced by $A$. We prove the following.
Theorem 1.1.
Let $A$ be a set of $n$ random points uniformly distributed in $[0,1]^{2}$, and let $n\rightarrow\infty$. Let $t(n)$ be an arbitrarily slowly growing function of $n$, with $\lim_{n\to\infty}t(n)=\infty$.
Then, with probability tending to $1$, the edge set of $K_{n}[A]$ can be covered by at most $O(n^{4/3}t(n))$ noncrossing paths, and, hence, by $O(n^{4/3}t(n))$ noncrossing matchings.
This is better than the $O(n^{3/2})$ bound of Araujo et al. [3], which holds for any point set in general position.
Note that a covering of the edge set of $K_{n}[A]$ by $O(n^{4/3}t(n))$ noncrossing matchings yields a decomposition (i.e., partition) of $K_{n}[A]$ into $O(n^{4/3}t(n))$ noncrossing matchings (via deletion of duplicate edges). However, the same is not true for coverings by paths.
For dense sets, the $O(n^{3/2})$ upper bound [3] for decompositions into plane matchings can be strengthened by requiring that the paths are monotone.
Theorem 1.2.
Let $A$ be an $\alpha$-dense point set in general position in the plane.
The edge set of $K_{n}[A]$ can be covered by $O(n^{3/2})$ monotone paths, and such a covering can be computed in $O(n^{2})$ time. Consequently, the edge set of $K_{n}[A]$ can be decomposed into $O(n^{3/2})$ monotone matchings.
From the opposite direction, we suspect that covering the edge set of the complete geometric graph of any dense point set requires a superlinear number of monotone paths.
Conjecture 1.3.
For some (possibly, for any) dense $n$-element point set $A$ in general position in the plane, covering the edge set of $K_{n}[A]$ requires a superlinear number of monotone paths.
In contrast to random or dense point sets, arbitrary point sets require a quadratic number of monotone paths in the worst-case.
Theorem 1.4.
Let $A$ be a set of $n$ points in general position in the plane.
Then the edge set of the complete geometric graph $K_{n}[A]$ can be covered by $n^{2}/6+O(n)$ monotone paths. On the other hand, there exist $n$-element point sets that require at least $n^{2}/15$ monotone paths.
The rest of this paper is organized as follows. In Sections 2, 3, and 4, we prove Theorems 2, 3, and 5, respectively. The last section contains some concluding remarks and open problems.
2 Proof of Theorem 1.1
Let $A=a_{1},\ldots,a_{n}$ be a sequence of $n$ points, randomly and uniformly distributed in $U=[0,1]^{2}$. Write $m={n\choose 2}$ and let $e_{1},\ldots,e_{m}$ be the edges of $K_{n}[A]$ listed in lexicographic order. Consider the following procedure for covering the edge set of $K_{n}[A]$ by monotone paths.
Let $\theta=cn^{-4/3}$ be a small angle, for a suitable small constant $c>0$, so that $N=\pi/\theta$ is an integer. Observe that sorting with respect to $\mathbf{v}$ and $-\mathbf{v}$ are equivalent. Let $\mathcal{V}$ be a set of $N$ unit vectors uniformly distributed in $[0,\pi)$ so that $(1,0)\in\mathcal{V}$. Observe that every monotone path naturally decomposes into two noncrossing matchings. For each vector $\mathbf{v}\in\mathcal{V}$, compute the $\mathbf{v}$-monotone spanning path $\xi$ induced by $A$ and output the two $\mathbf{v}$-monotone matchings whose union is $\xi$. Consequently, we have obtained $O(n^{4/3})$ monotone paths (or matchings). This forms Phase I of the covering.
We show that the expected number of edges that do not appear in these $O(n^{4/3})$ matchings is $O(n^{5/3})$. Further decompose the geometric graph $G=(V,E)$ on this edge set into $O(n^{4/3})$ noncrossing matchings (or paths), as outlined below. This forms Phase II of the covering. Consequently, the expected number of noncrossing matchings generated by the above procedure is $O(n^{4/3})$ in each of the two phases, hence $O(n^{4/3})$ overall.
Analysis.
We start with Phase I. Let $a,b\in A$ be any pair of points. Instead of working with monotone paths in many directions, we may assume for convenience that $ab$ is a vertical segment, and that the unit square has been rotated by some angle, correspondingly. We denote by $U_{\alpha}$ the square rotated by $\alpha$; we have ${\rm Area}(U_{\alpha})=1$ for any $\alpha$. Consider the geometric locus $R(a,b)$ of the points $p\in U_{\alpha}$ on lines making an angle of at most $\theta$ with $ab$ and which intersect $ab$, see Fig. 1.
Figure 1: Left: four lines making an angle of $\theta$ with $ab$ through $a$ and $b$. Center: the geometric locus (shaded) associated with a vertical segment. Right: a bounding rectangle for the locus associated with a long vertical segment.
Lemma 2.1.
The area of $R(a,b)$ is at most $4\theta$.
Proof 2.2.
Consider four lines incident to the $a$ and $b$ respectively, and making a (clockwise or counterclockwise) angle of $\theta$ with $ab$. The boundary of the locus is made by these four lines and two polygonal arcs on the boundary of the square. The locus is easily seen to be contained in an axis-parallel rectangle of width $2\sqrt{2}\tan\theta$ and height $\sqrt{2}$. As such, its area is bounded from above by $4\tan\theta$. Excluding the two isosceles triangles based on the left and right vertical sides yields an improved area bound of $\frac{3}{4}\cdot 4\tan\theta=3\tan\theta\leq 4\theta$ (here we assume that $n$ is sufficiently large).
Observe that if the region $R(a,b)$ is empty of points in $A\setminus\{a,b\}$, then the pair $a,b$ generates an edge in any $\mathbf{v}$-monotone spanning path induced by $A$, such that $\mathbf{v}\in\mathcal{V}$ makes an angle at most $\theta$ with the direction orthogonal to $ab$. In this case, $ab$ appears in part I of the decomposition. We quantify the probability of this event.
We say that pair $a,b$ is good if $ab$ is an edge of a $\mathbf{v}$-monotone spanning path considered in Phase I, with $\mathbf{v}$ making an angle at most $\theta$ with a line orthogonal to $ab$; and it is bad otherwise. By the principle of deferred decisions, see, e.g., [12], we may assume that $a,b$ are the first two points in the sequence of $n$ random draws when considering this particular $\mathbf{v}$. By using the independence of trials and Bernoulli’s inequality, we obtain:
${\rm Prob}\left(a,b\text{ is good}\right)$ $\geq\left(\frac{1-{\rm Area}(R(a,b)}{{\rm Area}(U_{\alpha})}\right)^{n-2}$
$>(1-{\rm Area}(R(a,b))^{n}\geq(1-4\theta)^{n}\geq 1-4n\theta$
$\geq 1-4cn^{-1/3}.$
It follows that
${\rm Prob}\left(a,b\text{ is bad}\right)\leq 4cn^{-1/3}.$ (1)
Recall that $e_{1},\ldots,e_{m}$ are the edges of $K_{n}[A]$ in a fixed (lexicographic) order. Let $Y_{i}$, $i=1,\ldots,m$ be the indicator random variable of the event that $e_{i}$ is a bad edge and let $Y=\sum_{i=1}^{m}Y_{i}$. By (1), the expected number of edges not appearing in monotone paths generated in Phase I is bounded from above as follows:
${\rm E}[Y]\leq m\cdot 4cn^{-1/3}\leq 2cn^{5/3}=:\mu.$
We use Markov’s inequality, see, e.g., [12, Thm. 3.1], to show that the bound we obtain on the number of monotone paths holds with probability tending to $1$ as $n\to\infty$.
${\rm Prob}\left(Y\geq\mu\cdot t(n)\right)\leq(t(n))^{-1}\to 0.$ (2)
We continue with Phase II. Let $G=(V,E)$, with $|E|=Y$, be the random geometric graph of bad edges surviving Phase I. For decomposing the edge set of $G$ into noncrossing matchings (or paths), we use a result of Tóth [19], according to which any geometric graph on $n$ points and with no $k+1$ disjoint edges has at most $2^{9}k^{2}n=O(k^{2}n)$ edges. Moreover, if the graph has more than $2^{9}k^{2}n$ edges, one can find $k+1$ pairwise disjoint edges, or a noncrossing path of $12k+$ 1 edges.
In our case, solving for $k$ in $n^{5/3}=nk^{2}$ yields $k=n^{1/3}$. Start with the graph $G$ on the $n$ points, and repeatedly remove a large set of pairwise disjoint edges, until the graph becomes empty. The process consists of at most $\lceil\log{(2cn^{5/3})}\rceil$ steps, $i=2,\ldots,\lceil\log{(2cn^{5/3})}\rceil+1$. In step $i$, the current graph has $m_{i}$ edges, where
$\frac{2cn^{5/3}}{2^{i}}<m_{i}\leq\frac{2cn^{5/3}}{2^{i-1}}.$
In step $i$, repeatedly find and remove $k_{i}+1$ disjoint edges (a matching), or a path of $22k_{i}+$ 2 edges, where $k_{i}\sim n^{1/3}2^{-i/2}$, until $m_{i}$ fails to satisfy the above inequality, and the process continues with step $i+1$. The number of matchings (or paths) removed in step $i$ is about
$\frac{n^{5/3}}{2^{i}}\cdot\frac{2^{i/2}}{n^{1/3}}=n^{4/3}\cdot 2^{-i/2},$
and so the total number of matchings (or paths) is at most
$n^{4/3}\sum_{i=2}^{\infty}2^{-i/2}=O(n^{4/3}).$
A final note: the paths (or matchings) that are generated by the procedure in [19] are not necessarily monotone. ∎
3 Proof of Theorem 1.2
Consider an $\alpha$-dense point set $A$. By scaling, we can assume that $A$ is contained in a unit square $U=[0,1]^{2}$, and the distance between any two points of $A$ is at least $\sqrt{2}\cdot\alpha n^{-1/2}$.
The covering procedure is similar to Phase I in the proof of Theorem 1.1, but with a twist. Let $\varphi=cn^{-1/2}$ be a small angle, for a suitable small constant $c>0$, so that $N=2\pi/\varphi$ is an integer. Let $\mathcal{V}$ be a set of $N$ unit vectors uniformly distributed around the origin so that $(1,0)\in\mathcal{V}$. For each vector $\mathbf{v}\in\mathcal{V}$, compute the $\mathbf{v}$-monotone spanning path $\xi_{v}$ induced by $A$. Let the path be $a_{1}a_{2}\ldots a_{n}$ in the order of monotonicity. Set
$\beta=\frac{8c+4\sqrt{2}\cdot\alpha}{\pi\alpha^{2}},$
and for each $1\leq j\leq i\leq\beta\sqrt{n}$, add the path $a_{j}a_{i+j}a_{2i+j}\ldots$ to the set $\mathcal{P}$ of covering paths. Note that for each vector $\mathbf{v}\in\mathcal{V}$, there are at most $1+2+\cdots+\beta\sqrt{n}<\beta^{2}n$ paths to be added to $\mathcal{P}$, all of which are monotone and therefore $\mathcal{P}$ contains at most $\frac{2\pi}{c}\sqrt{n}\cdot\beta^{2}n=\Theta(n^{3/2})$ monotone paths.
We claim that the paths in $\mathcal{P}$ cover the complete geometric graph on $A$. That is, for any pair of points $a,b\in A$, we show that the segment $ab$ appears in at least one path in $\mathcal{P}$. We construct the region $R(a,b)$ analogous to the one in the proof of Theorem 1.1 but with respect to $\varphi$ instead. By Lemma 2.1, its area is at most $4\varphi$. Assume that $R(a,b)$ contains $k$ points of $A$. The disks of radius $r=\frac{\sqrt{2}}{2}\cdot\alpha n^{-1/2}$ centered at these $k$ points are disjoint and they cannot exceed the region $R(a,b)$ extended by a margin of $r$, whose area is then at most $4(\varphi+r)$. The total area of the $k$ disks is at most $4(\varphi+r)$, and therefore
$k\pi r^{2}$ $\leq 4(\varphi+r),\text{ thus }$
$k\pi\frac{\alpha^{2}}{2n}$ $\leq 4\left(\frac{c}{\sqrt{n}}+\frac{\frac{\sqrt{2}}{2}\cdot\alpha}{\sqrt{n}}\right),\text{ or }$
$k$ $\leq\frac{8c+4\sqrt{2}\alpha}{\pi\alpha^{2}}\cdot\sqrt{n}=\beta\sqrt{n}.$
This means that in any $\mathbf{v}$-monotone spanning path induced by $A$, where $\mathbf{v}\in\mathcal{V}$ makes an angle of at most $\varphi$ with $ab$, there are at most $\beta\sqrt{n}$ points between $a,b$ and thus, the segment $ab$ appears in one of the monotone paths constructed from $\xi_{v}$.
Since every monotone path can be decomposed into two monotone matchings (the odd- and even-numbered edges in the path), the edge set of $K_{n}[A]$ can be covered by $\Theta(n^{3/2})$ monotone matchings.
For the analysis of the algorithm, we have $\Theta(\sqrt{n})$ vectors and for each vector $v$ we first sort the points according to their $v$ component in $O(n\log n)$ time. Finally, by tracing the points in the sorted array, each point may update at most $O(\sqrt{n})$ monotone paths corresponding to $v$. Overall, the running time of the algorithm is
$\Theta(\sqrt{n})\,\left[O(n\log n)+n\cdot O(\sqrt{n})\right]=O(n^{2}).$
4 Proof of Theorem 1.4
Lower bound. Assume that $n$ is divisible by $3$, i.e., $n=3k$. For convenience we denote our $n$-element point set by $P$. Partition the $n$ points into three groups $A,B,C$, of the same size, close to the vertices of a unit equilateral triangle, and each group consists of roughly equidistant points, but the distances are different; see Fig. 2. Let the diameters of the 3 groups of $k$ vertices be ${\rm diam}(C)\ll{\rm diam}(B)\ll{\rm diam}(A)\ll 1$.
The number of (undirected) inter-group edges, i.e., edges in
$E_{0}:=E(A\times B)\cup E(A\times C)\cup E(B\times C),$
is $33k^{2}=n^{2}/$ 3. Let $\mathcal{P}$ be a covering of $K_{n}[A]$ by monotone paths. We claim that every path $\xi\in\mathcal{P}$ contains at most five edges in $E_{0}$. Consequently, covering all the edges of the tripartite graph on $P=A\cup B\cup C$ requires at least $153k^{2}/5=n^{2}/$ 15 monotone paths.
Figure 2: Covering the edge set of a tripartite graph by monotone paths. The figure shows the edges in $E_{0}$ corresponding to a monotone path $\xi$; note that this edge set may be disconnected. A direction of monotonicity is drawn in blue color.
To verify the claim, observe that the directed edges in $\xi$ can be of six types:
  1. 1.
    $A\to B$, and the corresponding opposite orientation $B\to A$
  2. 2.
    $A\to C$, and the corresponding opposite orientation $C\to A$
  3. 3.
    $B\to C$, and the corresponding opposite orientation $C\to B$
Let us trace the edges of $\xi$ from one end of the path to the other, and record the sequence of groups in the order they are visited. (When two consecutive vertices belong to the same group, we do not repeat the corresponding symbol in this sequence.)
Due to the monotonicity of $\xi$, the resulting sequence satisfies the following conditions:
  1. (i)
    There is no cycle of the form $A\to B\to C\to A$.
  2. (ii)
    For any directed edge type $U$, let $U^{r}$ denote the type of an edge with opposite orientation. There are no two types, $U$ and $V$, such that $\xi$ has edges belonging to all four types, $U$, $U^{r}$, $V$, and $V^{r}$. This follows from the fact that there is no straight-line on which the orthogonal projection of four such edges would be disjoint.
  3. (iii)
    For any two groups, $X,Y\in\{A,B,C\}$, there are at most four edges between $X$ and $Y$ in $\xi$. Moreover, if ${\rm diam}(X)\gg{\rm diam}(Y)$ and there are four edges between $X$ and $Y$ in $\xi$, then $\xi$ starts or ends at a point in $X$.
This implies that there are no monotone paths consisting of more than five edges in $E_{0}$, and the only monotone paths covering five edges are of the following form:
  1. 1.
    $A\to B\to A\to B\to A\to C$ (shown in Fig. 2).
  2. 2.
    $A\to C\to A\to C\to A\to B$
  3. 3.
    $B\to C\to B\to C\to B\to A$
Upper bound. Since every 2-edge path is monotone, one can decompose the edge set of $K_{n}[A]$ into $\lceil\frac{1}{2}{n\choose 2}\rceil$ 2-edge paths. We can improve upon this bound by using $3$-edge paths. Since not every $3$-edge path is monotone, we need special $3$-edge paths that are monotone. A $3$-edge path is a zig-zag path if its two endpoints lie in opposite open halfplanes determined by the middle edge of the path. It is easy to see that any $3$-edge zig-zag path is monotone.
The key is a result by Wilson [21] about decomposition of the edge set of complete graphs into complete subgraphs of the same size. The case we are interested in is as follows.
Lemma 4.1 (Wilson 1975).
Let $n$ be a large enough positive integer. Then the edge set of $K_{n}$ can be decomposed into subgraphs of the form $K_{6}$, if and only if $n$ satisfies two divisibility conditions, namely $15|n-$ 1 and $\binom{6}{2}|\binom{n}{2}$.
Moreover, if $n$ does not satisfy the above divisibility conditions, then one can still pack $n^{2}/30-O(n)$ $K_{6}$’s into $K_{n}$, leaving out only linearly many edges.
Figure 3: Covering the edge set of the complete graph on six points by $3$-edge zig-zag paths.
Our approach is packing as many $K_{6}$ as possible into $K_{n}[A]$, then decomposing each $K_{6}$ into five $3$-edge zig-zag paths, and considering each of the remaining edges as a single path. This way we can guarantee a decomposition of the edge set of $K_{n}[A]$ into $\binom{n}{2}/\binom{6}{2}\times 5+O(n)=n^{2}/6+O(n)$ monotone paths.
It suffices to show that for a $6$-point set $A$, the edge set of $K_{6}[A]$ can be decomposed into five $3$-edge zig-zag paths. It is known that there are $16$ order types for such sets, see, e.g., [17], and it suffices to provide zig-zag path decompositions for each type, as we do in Fig. 3. ∎
A similar bipartite version of the lower bound construction with $n=2k$ points and $|E_{0}|=k^{2}$ (undirected) inter-group edges only yields a lower bound of $n^{2}/16$: If $\mathcal{P}$ is a covering of $K_{n}[A]$ by monotone paths, it has the property that that every path $\xi\in\mathcal{P}$ contains at most four edges in $E_{0}$. Consequently, covering all the edges of the bipartite graph requires at least $k^{2}/4=n^{2}/16$ monotone paths.
5 Concluding remarks
1. For $k\geq 2$, let $f(k,n)$ denote the maximum number of edges of a geometric graph on $n$ vertices that contains no $k$ pairwise disjoint edges. Early results of Erdős [7], Kupitz [11], and Alon and Erdős [2] show that $f(2,n)=n$ and $f(3,n)=O(n)$. The first polynomial upper bound for $f(k,n)$, for $k>3$, was established by Pach and Törőcsik [16], and was improved by Tóth and Valtr [20]; see also [14, Chap. 14]. The current best result, $f(k,n)=O(k^{2}n)$ is due to Tóth [19]. It was used in the proof of Theorem 1.1.
It is conjectured that $f(k,n)=O(kn)$ [5, Chap. 9.5]. If this conjecture is true, our methods would yield an $O(n\log{n})$ (unfortunately, still superlinear) upper bound on the number of monotone matchings sufficient for covering the edge set of $K_{n}[A]$, analogous to covering this set by noncrossing matchings in [3].
2. Obenaus and Orthaber [13] gave a negative answer to the question of whether every complete geometric graph on $n$ vertices ($n$ even) can be partitioned into $n/2$ spanning trees (see [4]). Furthermore, their negative answer extends to the weaker question of whether every complete geometric graph on $n$ vertices ($n$ even) can be partitioned into $n/2$ plane subgraphs. See also [1].
On the other hand, it is possible that every complete geometric graph on $n$ vertices can be partitioned into $n/2+o(n)$ plane subgraphs. Pach, Saghafian, and Schnider [15] proved that a complete convex geometric graph on $n$ vertices cannot be decomposed into fewer than $n-1$ plane star-forests.
3. Pinchasi and Yerushalmi [18] showed that, given any $n$-element point set $A$ in the plane where $n$ is odd, the edge set of $K_{n}[A]$ can be partitioned into $(n^{2}-1)/8$ convex polygons whose vertices belong to $A$. This bound is tight. For the case of even $n$, $n^{2}/8+n/4$ convex polygons suffice and $\lceil n^{2}/8\rceil$ are needed. Since every convex polygon can be decomposed into two monotone paths, we essentially re-obtain the upper bound on monotone paths that comes from covering by two-edge paths: if $n$ is odd, $(n^{2}-1)/4$ monotone paths suffice, whereas if $n$ is even, $n^{2}/4+n/2$ monotone paths suffice. These bounds are superseded by our Theorem 1.4.
4. Perhaps the first step toward solving Problem 1 would be to answer the following question.
Problem 5.1.
Does there exist a constant $c>1/2$ with the property that, for infinitely many values of $n$, there are $n$-element point sets $A$ in the plane such that any covering of the edge set of $K_{n}[A]$ by noncrossing paths uses at least $cn$ paths?
5. Another interesting class of paths is zig-zag paths. A path is called zig-zag if any three consecutive edges form a $3$-edge zig-zag path, as defined earlier. Note that this definition does not ensure that the path is monotone or even plane.
We claim that for any finite point set $A$ in general position in the plane, one can cover the edge set of $K_{n}[A]$ by plane Hamiltonian zig-zag paths. Equivalently, for any pair of points $a,b\in A$, the segment $ab$ appears in at least one such path. To see this, assume that the line through $a,b$ separates $A\setminus\{a,b\}$ into two disjoint parts $A_{1},A_{2}$. The edge $ab$ is an edge of the convex hull of $A_{i}\cup\{a,b\}$, for $i=1,2$. Starting with $ab$, we construct a plane zig-zag path that traverses all points in $A_{1}\cup\{a,b\}$, as follows. At the first step, let $b_{1}\in A_{1}$ be the point such that the angle $\angle abb_{1}$ is as small as possible. We continue this process by selecting $b_{i}$ at the $i$-th step such that the angle $\angle b_{i-2}b_{i-1}b_{i}$ is minimized among all remaining points in $A_{1}$, where $b_{0}=b$. It is not hard to verify that once all points in $A_{1}$ are included, the resulting path is a plane zig-zag path. Similarly, starting with $ba$ we can construct a plane zig-zag path that traverses all points in $A_{2}\cup\{a,b\}$. The union of these two paths is a plane zig-zag Hamiltonian path, as claimed.
While the edge set of any complete geometric graph can be covered by plane zig-zag (Hamiltonian) paths, the minimum number of (not necessarily Hamiltonian) plane zig-zag paths required for covering the edge set of $K_{n}[A]$ is not known.
Problem 5.2.
What is the smallest number $z=z(n)$ such that the edge set of every complete geometric graph on $n$ vertices can be covered by $z$ plane zig-zag paths?